# Improving the Efficiency of Self-Supervised Adversarial Training through Latent Clustering-Based Selection

### v1

@@ Abstract.
Compared with standard learning, adversarially robust learning is widely recognized to demand significantly more training examples. Recent works propose the use of self-supervised adversarial training (SSAT) with external or synthetically generated unlabeled data to enhance model robustness. However, SSAT requires a substantial amount of extra unlabeled data, significantly increasing memory usage and model training times. To address these challenges, we propose novel methods to strategically select a small subset of unlabeled data essential for SSAT and robustness improvement. Our selection prioritizes data points near the model’s decision boundary based on latent clustering-based techniques, efficiently identifying a critical subset of unlabeled data with a higher concentration of boundary-adjacent points. While focusing on near-boundary data, our methods are designed to maintain a balanced ratio between boundary and non-boundary data points to avoid overfitting. Our experiments on image benchmarks show that integrating our selection strategies into self-supervised adversarial training can largely reduce memory and computational requirements while achieving high model robustness. In particular, our latent clustering-based selection method with k-means is the most effective, achieving nearly identical test-time robust accuracies with $5$ to $10$ times less external or generated unlabeled data when applied to image benchmarks. Additionally, we validate the generalizability of our approach across various application scenarios, including a real-world medical dataset for COVID-19 chest X-ray classification. Our Implementations are available as open-source code at this url.
1. Introduction
@@
Over the past decade, it has been repeatedly confirmed that deep neural networks (DNNs) are highly vulnerable to adversarial perturbations (Szegedy et al., 2013). This phenomenon has raised serious concerns about the reliability of DNNs in safety-critical applications and has driven numerous research into designing methods to enhance model robustness (Goodfellow et al., 2014; Papernot et al., 2016; Buckman et al., 2018; Biggio and Roli, 2018). Among them, adversarial training is regarded as one of the most effective methods (Madry et al., 2017; Wang et al., 2019; Zhang et al., 2019). However, as stated by Schmidt et al. (Schmidt et al., 2018), learning a model that is resilient to adversarial perturbations requires a significantly larger dataset than standard learning. Recent studies have explored self-supervised techniques to expand the training set size of adversarial training algorithms by leveraging unlabeled external (Carmon et al., 2019) or generated data (Gowal et al., 2021; Sehwag et al., 2021; Wang et al., 2023). Despite alleviating the sample complexity barrier and producing models with improved robust accuracies, these methods typically utilize vast amounts of additional data, suggesting the requirement of much larger hardware to store those data and a much longer training time for algorithms like adversarial training to converge.
@@
Witnessing the challenges of additional memory and computational requirements, we investigate whether the significantly large amount of utilized additional data is inevitable for achieving state-of-the-art adversarial robustness. The ultimate goal of our work is to maximize the model robustness achieved by SSAT algorithms by using as few additional unlabeled data points as possible. Inspired by Zhang et al. (Zhang et al., 2020), which highlights the unequal importance of training examples, we argue that with limited model capacity, self-supervised adversarial learning should also focus on optimizing critical data samples near the model’s decision boundary. Consequently, we propose multiple data selection strategies, including a simple prediction confidence-based selection (PCS) scheme that prioritizes unlabeled data points in which the model’s prediction is most uncertain and more advanced latent clustering-based selection (LCS) schemes. Specifically, we propose two LCS approaches depending on the corresponding clustering technique it builds on: latent clustering-based selection with Gaussian mixture models (LCS-GMM) and latent clustering-based selection with k-means (LCS-KM). For each unlabeled data point, we compute the difference between the highest and second-highest posterior probabilities in the latent feature space for LCS-GMM, while we select vulnerable points by calculating the distance difference between the two nearest cluster centroids for LCS-KM. The set of unlabeled data with the smallest probability or distance difference is considered more critical and further incorporated into self-supervised adversarial training. Moreover, our selection strategies balance the ratio of boundary critical points and the remaining non-boundary points to avoid undesirable overfitting to the boundary data distribution.
By applying these targeted data reduction approaches, we streamline the self-supervised adversarial training process, significantly reducing the required additional unlabeled data while maintaining comparable model robustness against adversarial perturbations. Our work offers valuable insights into how to acquire additional data and employ SSAT in resource-constrained application domains. Below, we further summarize the main contributions of our work.
Contributions. We motivate and formalize the problem task of reducing the volume of unlabeled data while maintaining model robustness for SSAT (Section 3). To realize such a goal, we propose various data selection schemes to identify the most critical unlabeled data points, including a straightforward approach based on prediction confidence and two advanced methods based on clustering techniques of unlabeled data in the latent embedding space (Section 4). The proposed methods optimize the performance of SSAT by refining the model’s decision boundary in the input regions of high uncertainty by strategically prioritizing boundary-adjacent unlabeled data points. By focusing on critical unlabeled data points, our methods largely reduce the memory consumption and time complexities of self-supervised adversarial training algorithms. We conduct comprehensive experiments on two image benchmarks and a real-world medical dataset, demonstrating that our data selection schemes significantly reduce the memory consumption and the total running time of self-supervised adversarial training algorithms under various experimental settings (Section 5). Additionally, we explain the insights of the potential reasoning behind the superior efficacy of our LCS-KM method to other data selection schemes (Section 6.1) and provide general guidelines on how to employ our LCS methods for maximized robustness performance (Section 6.2). Finally, we examine the robustness of our methods by varying the adversarial training algorithms, perturbation size and metric (Section 6.3), as well as studying the role of the intermediate model under the regime with insufficient labeled data (Section 6.4), all confirming the usefulness and generalizability of the proposed data selection schemes for improving the efficiency of SSAT algorithms. We believe our work can help promote the development of more scalable, robust learning algorithms, which would be particularly beneficial for resource-constrained environments.
2. Related Work
In this section, we review the most relevant literature on adversarial examples, (self-supervised) adversarial training, and its variants. We also discuss the most representative techniques for improving the efficiency of standard deep learning, including dataset distillation and coreset selection.
2.1. Adversarially Robust Learning
Adversarial Examples. Adversarial examples are inputs crafted with small perturbations that are designed to mislead model predictions (Szegedy et al., 2013). The prevalence of adversarial examples with deep neural networks poses a critical challenge in modern machine learning, especially for security-critical or safety-critical applications. Earlier-proposed attacks like fast gradient sign method (FGSM) (Goodfellow et al., 2014) and projected gradient descent (PGD) (Madry et al., 2017) introduced gradient-based techniques to find adversarial examples using constrained optimization, while Carlini and Wagner proposed unconstrained CW attacks (Carlini and Wagner, 2017), which search for the minimal perturbation that causes misclassification. Recent attacks like Auto-PGD (Croce and Hein, 2020a) automate gradient steps for higher efficiency in finding worst-case perturbations, while AutoAttack (Croce et al., 2020) is an ensemble of diverse parameter-free attacks, making them more effective in evaluating model robustness.
Adversarial Training. On the defense side, adversarial training and its variants are the most popular learning methods for adversarial robustness. For instance, Madry et al. proposed to train DNNs on adversarially perturbed inputs using PGD attacks (Madry et al., 2017), while Zhang et al. introduced TRADES (Zhang et al., 2019), which balances the trade-off between adversarial robustness and standard accuracy. Nevertheless, adversarial training algorithms are often criticized for overfitting adversarially perturbed samples produced during training time and for their high computation costs, which hinder their deployment in real-world applications. In particular, Rice et al. illustrated through comprehensive experiments that the robust overfitting phenomenon is prevalent in adversarial training (Rice et al., 2020), where early stopping is shown to be effective in alleviating such an issue. Techniques, such as Free AT (Shafahi et al., 2019) and Fast AT (Wong et al., 2020), offer promising solutions to reduce the computation for adversarial training. Nevertheless, they suffer from decreased robustness performance.
Moreover, a line of research pointed out that treating all the data points equally in adversarial training is not optimal for robust learning since different examples can impact model robustness differently. For example, Zhang et al. argued that data points far from decision boundaries are inherently more secure and less likely to be influential in robust learning against adversarial perturbations (Zhang et al., 2020). As a result, these points should be given less weight during training to prevent the model from overfitting them, which could negatively affect its generalization on more vulnerable or unseen data. In a similar vein, Hua et al. advocated for a more targeted approach (Hua et al., 2021), recommending that PGD training be applied primarily to examples situated near the decision boundaries. This strategy focuses the model’s attention on the points that are most likely to enhance its robustness, as adversarial perturbations on boundary-adjacent data are more likely to result in meaningful improvements to the model’s resistance to attacks. Our work builds on these ideas by strategically identifying and prioritizing boundary-adjacent points from the unlabeled dataset to address the unique challenges associated with self-supervised adversarial training methods.
Self-Supervised Adversarial Training. Utilizing unlabeled data to improve model performance has become an active and rapidly advancing field of research (Chapelle et al., 2009). In the adversarial context, Schmidt et al. (Schmidt et al., 2018) highlighted a key observation: the sample complexity required for robust learning far exceeds that of standard learning. They showed that achieving adversarially robust generalization often necessitates a much larger dataset than traditional learning approaches, posing a challenge when labeled data is scarce or costly. Subsequently, a line of works proposed to boost model robustness by involving additional unlabeled data, often acquired from a similar but slightly different distribution, using self-supervised methods in adversarial training (Carmon et al., 2019; Alayrac et al., 2019; Najafi et al., 2019; Zhai et al., 2019). In particular, Carmon et al. proposed robust self-training methods to address the sample complexity issue (Carmon et al., 2019), which first trains an intermediate model using available labeled data and then leverages the model to generate pseudo-labels for unlabeled data. The pseudo-labeled samples, combined with the original labeled data, are subsequently used to train a final robust model. Concurrently, Najafi et al. proposed an optimization framework incorporating labeled and unlabeled data (Najafi et al., 2019), while providing formal guarantees on adversarial robustness.
In addition, leveraging synthetic data produced by state-of-the-art generative models to build robust models has also been explored extensively in recent literature (Gowal et al., 2021; Sehwag et al., 2021; Wang et al., 2023). For instance, Gowal et al. employed various unconditional generative models (Gowal et al., 2021), including variational autoencoders (VAEs), generative adversarial networks (GANs), and the more advanced denoising diffusion probabilistic models (DDPMs), to produce a large synthetic dataset to be incorporated in adversarial training, which improves robustness further. While all the aforementioned SSAT methods enrich the training dataset, they require considerable additional unlabeled data to ensure effective robustness enhancement. In this work, we explore data selection schemes to reduce the amount of unlabeled or generated data that are both efficient and effective for self-supervised adversarial training to achieve good robustness performance.
2.2. Data-Efficient Deep Learning
Dataset Distillation. Dataset distillation compresses large datasets into compact, synthetic datasets by matching gradients or optimizing specific loss functions, allowing models to achieve comparable performance with significantly fewer data points. For instance, Wang et al. proposed a foundational approach to dataset distillation (Wang et al., 2018), proposing methods to condense large datasets into smaller, synthetic subsets while preserving model performance. Later, Zhao et al. introduced the concept of gradient matching (Zhao et al., 2020), where distilled data points are optimized to match the gradients of the original dataset. This technique largely improves the distillation performance, enabling models to learn more effectively from reduced data. Additionally, Sucholutsky and Schonlau proposed soft-label dataset distillation, which uses probabilistic labels for distilled data (Sucholutsky and Schonlau, 2021), facilitating the training of complex models with minimal data while maintaining strong performance. Despite its success in standard deep learning, dataset distillation is often impractical for robust learning due to its high computational demands. Moreover, it assumes that all the examples in the entire dataset are equally important for distillation, ignoring the fact that different data points may have varying influences on model performance.
Coreset Selection. On the other hand, coreset selection (Sener and Savarese, 2017; Kim and Shin, 2022) aims to identify a small, representative data subset that retains the model performance when trained on this reduced dataset. These approaches often rely on diversity-based metrics or optimization techniques to ensure that the selected subset preserves sufficient information for effective learning. For example, Kaushal et al. utilized diverse models to select training data subsets to reduce labeling efforts (Kaushal et al., 2019), while Xia et al. introduced the concept of “moderate coreset” (Xia et al., 2022), focusing on data points with scores near the median to construct subsets that generalize well across different scenarios based on score distributions. Similarly, Mirzasoleiman et al. proposed selecting a weighted subset of training data that approximates the full gradient by maximizing a submodular function (Mirzasoleiman et al., 2020).
In the adversarial context, Dolatabadi et al. proposed adversarial coreset selection (Dolatabadi et al., 2023), offering a task-specific solution to enhance the computational efficiency of adversarial training. In particular, the method constructs a compact data subset by minimizing the gradient approximation error. The constructed subset is then utilized for adversarial training, with selection taking place at regular intervals. While effective in reducing computational overhead for vanilla adversarial training, the growing reliance on large volumes of generated or external data—such as in self-supervised adversarial training—introduces new challenges. These approaches often expand the dataset size to improve robust accuracy, eventually increasing the total training time. In such settings, repeatedly selecting data during training may become computationally infeasible. In summary, our work complements the above-mentioned methods for efficient deep learning, focusing on developing more data selection schemes without compromising robustness or scalability, particularly for SSAT when large-scale unlabeled data are involved.
3. Improving Data Efficiency for SSAT
In this section, we first introduce the necessary notations and definitions for readers to understand self-supervised adversarial training (Section 3.1) and then motivate and formulate the problem task of enhancing the data efficiency of SSAT algorithms (Section 3.3).
3.1. Preliminaries on SSAT
@@
We work with self-supervised adversarial training algorithms, initially introduced to the field by Carmon et al. (Carmon et al., 2019). Let $\mathcal{X}\subseteq\mathbb{R}^{n}$ be a $n$-dimensional input space, $\mathcal{Y}$ be the output space of class labels, and $D_{l}$ be the underlying labeled distribution supported on $\mathcal{X}\times\mathcal{Y}$ that we aim to learn an adversarially robust classifier. Let $\mathcal{S}_{l}$ be the training set with examples i.i.d. drawn from $D_{l}$ and $\mathcal{S}_{u}$ be a set of inputs sampled from some unlabeled distribution $D_{u}$ supported on $\mathcal{X}$. In the adversarial machine learning literature, $D_{u}$ is typically considered to be similar but not identical to the marginal input distribution of $D_{l}$, and $|\mathcal{S}_{u}|$ is set to be much larger than $|\mathcal{S}_{l}|$, since unlabeled data are easier to acquire than well-annotated labeled data. Throughout the paper, we use ||$\cdot$|| ⋅ | to denote the cardinality of the corresponding set. For instance, when $\mathcal{S}_{l}$ corresponds to the $50$K labeled CIFAR-10 training images, Carmon et al. (Carmon et al., 2019) selected $500$K “most CIFAR-10-like but non-identical” images from the whole $80$M Tiny ImageNet data as the unlabeled dataset $\mathcal{S}_{u}$ to be used by SSAT.
@@
To be more specific, SSAT first standardly trains an intermediate model $f_{\hat{\theta}}$ on $\mathcal{S}_{l}$, known as the pseudo labeling function, and then assigns pseudo-labels to each unlabeled data point $\bm{x}\in\mathcal{S}_{u}$. More rigorously, the intermediate model’s weight parameters $\hat{\theta}$ are learned by minimizing the following objective function:
(1) $\hat{\theta}=\operatorname*{argmin}_{\theta}\ \frac{1}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\mathcal{L}(\theta,\bm{x},y),$
@@
where $\mathcal{L}(\cdot)$ denotes the standard loss, such as cross-entropy, that measures the discrepancy between model prediction and the class label. Finally, SSAT trains a model, denoted as $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{S}_{u},\gamma)$, on both labeled dataset $\mathcal{S}_{l}$ and unlabeled dataset $\mathcal{S}_{u}$ but paired with pseudo labels by minimizing the following adversarial loss:
@@
(2) $\min_{\theta}\ \frac{1}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\mathcal{L}_{\mathrm{adv}}(\theta,\bm{x},y)+\frac{1}{\gamma|\mathcal{S}_{u}|}\sum_{\bm{x}\in\mathcal{S}_{u}}\mathcal{L}_{\mathrm{adv}}\left(\theta,\bm{x},f_{\hat{\theta}}(\bm{x})\right),$
@@
where $\gamma\geq 0$ controls the contributions between labeled and unlabeled data. In prior literature (Gowal et al., 2021), a typical approach is to construct each training batch with varying ratios of labeled and unlabeled data (corresponding to different values of $\alpha$) but fix the total batch size to optimize for the best SSAT performance. For simplicity, we term $\gamma$ as the per-batch ratio hyperparameter in this paper. When $\frac{1}{\gamma}$ is set as 00, Equation 2 reduces to the training objective of vanilla adversarial training (AT) (Madry et al., 2017). In Equation 2, $\mathcal{L}_{\mathrm{adv}}(\cdot)$ stands for the adversarial loss function: for any $(\bm{x},y)\in\mathcal{S}_{l}$ and $\epsilon>0$,
(3) $\mathcal{L}_{\mathrm{adv}}(\theta,\bm{x},y)=\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\mathcal{L}(\theta,\bm{x}+\bm{\delta},y),$
@@
where $\mathcal{B}_{\epsilon}(\bm{0})$ denotes the ball centered at $\bm{0}$ with radius $\epsilon$ measured in some distance metric, such as $\ell_{p}$-norm. Aligned with prior literature on adversarial training (Madry et al., 2017), we adopt multi-step PGD to obtain an approximated solution to the inner maximization problem specified by Equation 3 during self-supervised adversarial training.
3.2. Limitations of Existing SSAT Methods

### v2

# Efficient Semi-Supervised Adversarial Training via Latent Clustering-Based Data Reduction

@@ Abstract
Achieving high model robustness under adversarial settings is widely recognized as demanding considerable training samples. Recent works propose semi-supervised adversarial training (SSAT) methods with external unlabeled or synthetically generated data, which are the current state-of-the-art. However, SSAT requires substantial extra data to attain high robustness, resulting in prolonged training time and increased memory usage. In this paper, we propose unlabeled data reduction strategies to improve the efficiency of SSAT. Specifically, we design novel latent clustering-based techniques to select or generate a small critical subset of data samples near the model’s decision boundary. While focusing on boundary-adjacent points, our methods maintain a balanced ratio between boundary and non-boundary data points to avoid overfitting. Comprehensive experiments on benchmark datasets demonstrate that our methods can significantly reduce SSAT’s data requirement and computation costs while preserving its strong robustness advantages. In particular, our latent-space selection scheme based on k-means clustering and our guided DDPM fine-tuning approach with LCG-KM are the most effective, achieving nearly identical robust accuracies with $5\times$ to $10\times$ less unlabeled data and approximately $4\times$ less total runtime.
I Introduction
@@
Deep neural networks (DNNs) are highly vulnerable to adversarially perturbed inputs [1, 2]. This phenomenon has raised serious concerns about the reliability of DNNs in safety-critical applications and has driven numerous research into designing methods to enhance model robustness [2, 3, 4, 5]. Among them, adversarial-training-based algorithms are most popular [7, 8, 9]. As illustrated by Schmidt et al. [10], adversarially robust learning is inherently challenging, requiring a much larger sample complexity than standard learning. To address this, semi-supervised adversarial training (SSAT) has been proposed to enlarge the training set size with either external unlabeled data [11, 12] or generated synthetic data [13, 14, 15]. Despite alleviating the sample complexity barrier and producing models with improved robustness performance, these SSAT algorithms typically utilize vast amounts of additional data, suggesting the requirement of much larger hardware to store the extra data and a much longer training time for algorithms like adversarial training to converge (see Appendix A for detailed discussions on related work).
@@
Motivated by the above observations, we propose to study whether the large amount of extra unlabeled data is inevitable for SSAT to achieve high robust accuracy. The ultimate goal is to maximize model robustness while utilizing as few additional data samples as possible. Inspired by Zhang et al. [16], which highlights the unequal importance of training examples, we argue that with limited model capacity, SSAT should also focus on critical points near the model’s decision boundary. Specifically, we introduce three data selection strategies, including PCS, a simple approach that prioritizes unlabeled data points with low prediction confidence, and two advanced latent clustering-based selection schemes: LCS-KM and LCS-GMM, where we identify the most uncertain data in the model’s latent embedding space based on the corresponding clustering technique. The selected data points are then incorporated into the SSAT pipeline as the unlabeled dataset. In addition, we propose novel DDPM fine-tuning approaches with guidance loss inspired by our data selection schemes, including PCG, LCG-KM, and LCG-GMM, to further reduce the computation overhead of SSAT algorithms that require the pre-generation of synthetic data using diffusion models. Our data reduction schemes balance the ratio of boundary critical points and the remaining non-boundary points to avoid potential overfitting to the boundary distribution. By applying the aforementioned data reduction schemes, the high unlabeled data and computation requirements of SSAT can be significantly alleviated while largely preserving its robustness performance.
Contributions. We motivate and formalize the problem tasks of reducing the volume of unlabeled data while maintaining model robustness for SSAT (Section II). To realize this goal, we propose to optimize the performance of SSAT by refining the model’s decision boundary in the regions of high uncertainty by strategically prioritizing boundary-adjacent unlabeled data points (Section III-A). Specifically, we propose various data selection schemes to identify boundary-adjacent points (Section III-B), including a simple method based on prediction confidence and two advanced selection techniques by clustering the unlabeled or generated data in the latent embedding space. Additionally, we introduce three guided DDPM fine-tuning approaches (Section III-C), aiming to generate the subset of desirable data directly, thus avoiding the computation overhead of pre-generating the whole extra dataset.
Comprehensive experiments on image benchmarks demonstrate that our data selection schemes, especially for the best-performing LCS-KM approach, significantly improve the data and computational efficiency of baseline SSAT methods, while achieving comparable robust accuracies (Section IV-A). For instance, utilizing $10\%$ extra unlabeled data identified by LCS-KM in SSAT can achieve similar robustness performance on SVHN compared to using the entire dataset, while reducing the total runtime to approximately $25\%$. By directly generating the exact amount of boundary-adjacent data points, our guided DDPM fine-tuning approach, such as LCG-KM, can reduce the computation overhead while maintaining robustness (Section IV-C). Moreover, we demonstrate the generalizability of our data reduction schemes to a real-world COVID-related medical application (Section V-A) and visualize the distributions of selected data points to further illustrate why our LCS-KM method always performs the best (Section V-B). Finally, we provide general guidelines on how to choose the key hyperparameters for maximizing the SSAT performance (Section V-C), and conclude with more discussions on related literature, limitations of our study, and future research directions (Section VI). Our work reveals the importance of prioritizing boundary-adjacent points for improving SSAT’s efficiency and highlights the superiority of latent-based k-means clustering, potentially assisting the development of more efficient and scalable robust learning algorithms for real-world applications.
II Inefficiency of Existing SSAT Methods
II-A Semi-Supervised Adversarial Training
@@
We work with adversarially robust classification tasks and focus on semi-supervised adversarial training (SSAT), which was initially introduced to the field by Carmon et al. [11]. Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ be the input space and $\mathcal{Y}$ be the output space of class labels. Assume $D_{l}$ is the underlying labeled data distribution over $\mathcal{X}\times\mathcal{Y}$ that we aim to learn adversarially robust classifiers. Let $\mathcal{S}_{l}$ be the training dataset with examples independently and identically drawn from $D_{l}$, and $\mathcal{S}_{u}$ be a collection of inputs sampled from an unlabeled data distribution $D_{u}$ defined over $\mathcal{X}$. In adversarial ML literature, $D_{u}$ is typically considered to be similar but not identical to the marginal input distribution of $D_{l}$, and the set size of $\mathcal{S}_{u}$, denoted as $|\mathcal{S}_{u}|$, to be much larger than $|\mathcal{S}_{l}|$, since unlabeled data are much easier to acquire than well-annotated labeled data. For instance, when $\mathcal{S}_{l}$ corresponds to the 5050K labeled CIFAR-10 training images, $\mathcal{S}_{u}$ is the 500500K “most CIFAR-10-like but non-identical” images selected from the whole 8080M Tiny ImageNet dataset [11].
@@
Specifically, SSAT usually first trains an intermediate model $f_{\hat{\theta}}$ on the labeled training set $\mathcal{S}_{l}$, known as the pseudo labeler, and then assigns pseudo-labels to each unlabeled data point $\bm{x}\in\mathcal{S}_{u}$. The weight parameters of the intermediate model are optimized by minimizing the following objective function:
$\hat{\theta}=\operatorname*{argmin}_{\theta}\ \frac{1}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\ell(\theta,\bm{x},y),$ (1)
@@
where $\ell(\cdot)$ stands for the individual standard loss (e.g., cross-entropy), capturing the discrepancy between model predictions and the ground-truth labels. Finally, SSAT adversarially trains a classification model, denoted as $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{S}_{u},\gamma)$, on both labeled data $\mathcal{S}_{l}$ and unlabeled data $\mathcal{S}_{u}$ but paired with pseudo labels by minimizing the following adversarial loss:
@@
$\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{S}_{u},$ $\gamma)=\operatorname*{argmin}_{\theta}\quad\frac{\gamma}{|\mathcal{S}_{l}|}\sum_{(\bm{x},y)\in\mathcal{S}_{l}}\ell_{\mathrm{adv}}(\theta,\bm{x},y)$
$+\frac{1-\gamma}{|\mathcal{S}_{u}|}\sum_{\bm{x}\in\mathcal{S}_{u}}\ell_{\mathrm{adv}}\!\left(\theta,\bm{x},f_{\hat{\theta}}(\bm{x})\right),$ (2)
@@
where $\gamma\geq 0$ is a hyperparameter that controls the trade-off between labeled and unlabeled data. A typical approach used in prior work is to construct each training batch with varying ratios of labeled and unlabeled data (corresponding to different values of $\gamma$) but fix the batch size to optimize for the best SSAT performance [13]. For simplicity, we term $\gamma$ as the per-batch ratio. When $\gamma$ is set as $1$, Equation II-A reduces to the optimization objective of vanilla adversarial training (AT) [7]. In Equation II-A, $\ell_{\mathrm{adv}}(\cdot)$ denotes the individual adversarial loss. Specifically, for any $(\bm{x},y)\in\mathcal{S}_{l}$ and strength $\epsilon>0$,
$\ell_{\mathrm{adv}}(\theta,\bm{x},y)=\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\ell(\theta,\bm{x}+\bm{\delta},y),$ (3)
@@
where $\mathcal{B}_{\epsilon}(\bm{0})$ denotes the perturbation ball centered at $\bm{0}$ with radius $\epsilon$ in some distance metric such as $\ell_{p}$-norm. Aligned with prior literature [7], we adopt multi-step PGD to approximately solve for the optimization problem in Equation 3.
II-B Limitations of Prior Work